OpenStudy (anonymous):

find all positive integers n for which \[n^2+n+7\] is a complete square

OpenStudy (kinggeorge):

I believe, the only solution will be \(n=1,6\). Give me a minute to type up why.

OpenStudy (kinggeorge):

First, notice that for it to be a positive square, we require the following relation to be true: \[n^2+n+7\geq n^2+2n+1\]Since \(n^2+2n+1=(n+1)^2\). If we solve that relation, we find that \(n\leq 6\). Use trial and error on \(n=1,2,3,4,5,6\) and you find \(n=1,6\) work.

OpenStudy (anonymous):

\[n^2< n^2+n+7<(n+3)^2 \\ n^2+n+7=(n+1)^2 \ \ or \ \ (n+2)^2 \\ n=6,1\]

OpenStudy (anonymous):

@KingGeorge nice work

OpenStudy (kinggeorge):

Thanks.